OpenStudy (anonymous):

Given a^(3-x) b^(5x) =a^(x+5) b^(3x) then xlogb/a=?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@traile

OpenStudy (anonymous):

i am afraid i do not understand the question

OpenStudy (anonymous):

@johnnyappleseed

OpenStudy (kropot72):

I got \[x \ln \frac{b}{a}=\ln a\]

OpenStudy (anonymous):

a^(3-x) b^(5x) =a^(x+5) b^(3x) a^(3-x-x-5) = b^(3x-5x) a^(-2x-2)=b^(-2x) a^(-2x) a^(-2) = b^(-2x) a^(-2) = (b/a)^(-2x) log(a^(-2))=log((b/a)^(-2x)) (-2)loga = -2x log(b/a) loga =x(log(b /a))

OpenStudy (anonymous):

a^(3-x)b^5x= a^(x+5) b^3x a^(3-x-x-5 )= b^(3x-5x) a^-2(x+ 1) =b^-2x then take log on both side -2(x+1)loga=-2xlogb then divide it by neg. two (x+1)loga=xlogb logb/loga=(x+1)/x logb/loga= 1 + (1/x) then subtract 1 to the other side 1/x = (logb- loga)/loga x log(a/b) Then you will get the answer

OpenStudy (anonymous):

@CoCoTsoi plzz help i didnt understand ..

OpenStudy (anonymous):

which step? :)

OpenStudy (anonymous):

a^(3-x-x-5) = b^(3x-5x)

OpenStudy (anonymous):

take a^2 / a^4 as example. (a)(a)/ ((a)(a)(a)(a)) = 1/(a^2) = a^(-2) = a^(2-4) Similarly, a^(3-x) b^(5x) =a^(x+5) b^(3x) a^(3-x) / a^(x+5) = b^(3x) / b^(5x) a^(3-x-x-5) = b^(3x-5x)

OpenStudy (anonymous):

ok @CoCoTsoi how a^(-2) = (b/a)^(-2x)??

OpenStudy (anonymous):

a^(-2x) a^(-2) = b^(-2x) a^(-2x) a^(-2) / (a^(-2x))= b^(-2x) / (a^(-2x)) a^(-2) = (b/a)^(-2x)

OpenStudy (anonymous):